1. Suppose a consumer's income is Birr 10.000. The consumer uses his entire income to buy me goods
and X If the price of X1 is Bin 100 and the price of X2 is Birr 120 per unit
A. Show the effect of fall in income to Birr 8.000 on the vertical intercept, horizontal intercept and
slope of the budget line
B. Show the effect if the price of good X1 decreases to 60 birr per unit as price of X2 holds constant
C.Show the effect if the price of good X2 decreases to Wire 100 bir per unit as price of good
remains constant
D. Show the effect on the budget line if prices of both commodities changed proportionately
1
Expert's answer
2021-10-17T16:49:27-0400

a) The budget line will shift inward since the consumer cannot purchase more products as before.

Original consumer budget constraint: "100X1 + 120X2 \u2264 10000."

New consumer budget Constraint: "100X1 + 120X2 \u2264 8000."

The new vertical intercept of the budget line: "\frac{8000}{100} = 80"

The new horizontal intercept of the budget line: "\frac{8000}{120} = 67"

The new slope of the budget line will be: "\frac{-80}{67}"


b)The consumer budget constraint will be:

"60X1 + 120X2 \u2264 10000."

The new vertical intercept: "\frac{10000}{60} = 167"

The new horizontal intercept: "\frac{10000}{60} = 83"

The new slope of the budget line= "\frac{-167}{83}"


c) The consumer budget constraint will be:

"100X1 + 100X2 \u2264 10000."

The new vertical intercept: "\frac{10000}{100} = 100"

The new horizontal intercept of the budget line: "\frac{10000}{100} = 100"

The new slope of the budget line= "-1"


d)When the goods prices of the goods are changed proportionally the units that can be purchased will also change proportionally.